# On the Limit of Language Models as Planning Formalizers

### v2

@@ Abstract
Large Language Models have been shown to fail to create executable and verifiable plans in grounded environments. An emerging line of work shows success in using LLM as a formalizer to generate a formal representation (e.g., PDDL) of the planning domain, which can be deterministically solved to find a plan. We systematically evaluate this methodology while bridging some major gaps. While previous work only generates a partial PDDL representation given templated and thus unrealistic environment descriptions, we generate the complete representation given descriptions of various naturalness levels. Among an array of observations critical to improve LLMs’ formal planning ability, we note that large enough models can effectively formalize descriptions as PDDL, outperforming those directly generating plans, while being robust to lexical perturbation. As the descriptions become more natural-sounding, we observe a decrease in performance and provide detailed error analysis. [^1] Our code and data can be found at https://github.com/CassieHuang22/llm-as-pddl-formalizer.
@@ 1 Introduction
Large language models (LLMs) can make informal plans, such as suggesting ideas for parties or giving general advice on immigration. However, most users, let alone automated agents like robots, would not be able to actually execute those plans step-by-step to fruition – either to organize parties or acquire visas – without significant prior knowledge or external help. This inability to make executable plans lies in LLMs’ inability of grounding and formal reasoning Liu et al. (2023b); Pan et al. (2023); Zhang et al. (2023). Cutting-edge research in the community has evaluated LLMs’ ability to make formal plans in grounded environments, such as textual simulations, where all objects and actions represent actualities in the real world. Therefore, any resulting plan that formally involves those objects and actions would be executable and verifiable by nature. Although formal planning has been desirable in the history of AI Weld (1999), recent work found that even state-of-the-art LLMs are unable to generate formal plans Silver et al. (2024); Valmeekam et al. (2024); Stechly et al. (2024).
@@
Instead of using the LLM as a planner to generate the plan directly, an alternative line of work uses the LLM as a formalizer. Here, the LLM generates a formal representation of a planning domain, for example in the planning domain definition language (PDDL), based on some natural language descriptions of the environment. This representation can then be fed into a solver to find the plan deterministically (see Figure 1). Previous work achieved great success by showing that LLM-as-formalizer greatly outperforms LLM-as-planner in various domains Lyu et al. (2023); Xie et al. (2023); Liu et al. (2023a); Zhang et al. (2024a); Zuo et al. (2024); Zhang et al. (2024c); Zhu et al. (2024), as LLMs are more capable of information extraction than formal reasoning Zhang et al. (2024b). However, the above work has two major shortcomings. First, to simplify the task and evaluation, most have only attempted to generate a partial PDDL representation while assuming the rest is provided, often unrealistic in real life. Second, the language used to describe the environments is often artificially templated and structured, leading to potential overestimation of models’ ability.
This paper explores the limit of LLM-as-formalizer devoid of the above two simplifications. We use LLMs to generate the entirety of a PDDL representation, including the domain file and the problem file, given a natural-sounding description of the environment and the task (see Figure 1). On 4 widely used planning simulations from the International Planning Competition, we benchmark a suite of LLMs on generating PDDL that is both solvable and correct. As the descriptions in these datasets are templated, we also construct model-generated, human-validated descriptions that are natural-sounding to different levels.
Our work is the first to systematically analyze state-of-the-art LLMs’ ability of the trending methodology of LLM-as-formalizer on the highly challenging task of formal planning. We put forward an array of observations that will benefit future efforts. Discussed in detail in Section 5, our key findings are as follows.
  • On various planning simulations, closed-source models like GPT can decently generate entire PDDL, while open-source models like Llama up to 405B cannot.
  • When feasible, LLM-as-formalizer greatly outperforms LLM-as-planner.
  • As the environment descriptions sound more human-like, the models are more challenged.
  • The performance of LLM-as-formalizer is robust to lexical perturbation, while that of LLM-as-planner is not.
  • Open-source models succumb to syntax errors unlike GPT models, while semantic errors are common for both types of models.
@@ 2 Task: Formal Planning with PDDL
Formal planning (or classical planning) with PDDL involves a domain file ($\mathbb{DF}$) and problem file ($\mathbb{PF}$) (Figure 1). $\mathbb{DF}$ describes general properties in a planning domain that holds true across problems, while $\mathbb{PF}$ describes specific configurations of each problem instance. Concretely, the $\mathbb{DF}$ defines all available actions (and their parameters and pre- and post-conditions) for a state-based environment as well as predicates that represent the properties of object types. The $\mathbb{PF}$ defines the involved objects, the initial states, and the goal states. These two files are then given to a deterministic planner which will algorithmically search for a plan. Such a plan is a series of executable, instantiated actions that sequentially transition the world states from initial to goal. In the AI community, classical planning has been deemed an effective approach to solve real-world users’ problems, as the process is precise, explainable, verifiable, and deterministic.
However, formal planning demands a well-crafted pair of $\mathbb{DF}$ and $\mathbb{PF}$. In a real-world planning scenario, an average user would not describe their environment and problem with PDDL, but more likely with a textual description of the domain ($\mathbb{DD}$) and the problem ($\mathbb{PD}$), which can be specific or loose. Hence, we focus on the textual flavor of formal planning: given $\mathbb{DD}$ and $\mathbb{PD}$, the model outputs a successful plan with regard to the $\mathbb{DF}$ and $\mathbb{PF}$ that are withheld from the model.
@@ 3 Methodology: LLM-as-Formalizer
To tackle the task above, recent work involving LLMs diverged into two methodologies. The first, LLM-as-planner, directly uses LLMs to generate a plan based on the $\mathbb{DD}$ and a $\mathbb{PD}$. The second, LLM-as-formalizer, uses LLMs to recover the $\mathbb{DF}$ and $\mathbb{PF}$, before using a deterministic planner to arrive at the plan (Figure 2). Our work will focus on the second while using the first as a baseline. LLM-as-formalizer is in essence neurosymbolic, where LLMs help define the structured representation that is otherwise prohibitively costly to annotate and brittle to generalize. Existing works in this line demonstrated success but only generated a partial PDDL representation, while assuming the rest, including $\mathbb{PF}$ goals Lyu et al. (2023); Xie et al. (2023), the $\mathbb{PF}$ Liu et al. (2023a); Zhang et al. (2024a); Zuo et al. (2024), the action semantics in the $\mathbb{DF}$ Zhang et al. (2024c); Zhu et al. (2024), and the domain file Wong et al. (2023); Guan et al. (2023). While this simplifies the task and evaluation, the assumption of provided PDDL components is often unrealistic. Therefore, we bridge this gap by asking the LLM to predict the entire PDDL – both the $\mathbb{DF}$ and $\mathbb{PF}$ [^2] It is however minimally necessary to provide the action space, the identifiers and parameters of the actions in $\mathbb{DF}$, so the agent knows what actions are possible.
@@ 4 Evaluation: Datasets, Metrics, Models
To evaluate both approaches above, we work with fully-observed textual environments. Here, the provided $\mathbb{DD}$ and $\mathbb{PF}$ contain all necessary information for the model to make a complete plan.
@@ 4.1 Datasets
We consider four simulated planning domains, BlocksWorld, Logistics, Barman from the International Planning Competition IPC (1998), and MysteryBlocksWorld Valmeekam et al. (2024).
BlocksWorld, also used in Liu et al. (2023a), is a domain to rearrange stacks of blocks on a table using a robotic arm. There is 1 type of entities, 5 predicates, and 4 actions.
Mystery BlocksWorld obfuscates the original BlocksWorld domain by replacing all the names of the types, predicates, actions, and objects with nonsensical words, akin to a wug test Berko (1958). This dataset as an control group can effectively test whether models create plans via lexical pattern-matching and memorization.
@@
Barman, also used in Zhu et al. (2024), is a domain to to create cocktails from ingredients using different containers and two robotic arms. In this domain, there are 7 types of entities, 13 predicates, and 12 actions.
Each dataset comes with ground-truth PDDL describing domains ($\mathbb{DF}$) and problems ($\mathbb{PF}$). The input to the model is a natural language description of the domain ($\mathbb{DD}$) and the problem ($\mathbb{DD}$). The output of the model is a plan, namely a sequence of instantiated actions defined in $\mathbb{DF}$. For each of these datasets, the natural language description $\mathbb{DD}$s and $\mathbb{PD}$s were created in 3 different levels of naturalness.
Heavily Templated. For BlocksWorld, Logistics and Barman, the heavily templated $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as MysteryBlocksworld Valmeekam et al. (2024). This description is almost a word-by-word translation of PDDL. For example, for the ‘pick-up’ action in BlocksWorld, the ground-truth PDDL $\mathbb{DF}$ would be the following:
@@
From an application point of view, spelling out all preconditions and effects in terms of the predicates is paradoxical, as it assumes the user already have the algorithmic awareness of PDDL.
Moderately Templated. The $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as the original BlocksWorld dataset, following Valmeekam et al. (2024). For example. the Moderately Templated description of the ‘pick-up’ action is:
@@
While more natural-sounding than the Heavily Templated version, the description still explicitly discusses the preconditions and effects as well as predicates like ‘clear’. Moderately Templated datasets are available only for BlocksWorld and Logistics due to the size of Barman.
@@
To generate $\mathbb{DD}$, we ask GPT-4o with high temperature to generate and paraphrase a seed annotated $\mathbb{DD}$, and then manually verify the correctness by making sure it lists the correct predicates, preconditions and effects, which are not unique. We next verify the naturalness of the generated text by making there were variations in language throughout all descriptions generated, but was not giving out unnecessary information.
To generate $\mathbb{PD}$, we provide the model with a symbolic configuration that contains the number blocks, the initial stack configuration and the goal stack configuration. The model then ‘humanize’ the problem by making it sound natural, given a couple of seed exemplars. We manually verify the correctness of the dataset of the non-templated problems by hand by comparing them against the problem configurations. We then verify the naturalness of the $\mathbb{PD}$ by making sure there is variation but no ambiguity in its language.
Figure 3: Performance of both usages of LLMs on Heavily Templated BlocksWorld-100, Logistics-100, and MysteryBlocksWorld-100. Detailed results are shown in Appendix C.
@@
The above example of the Natural description no longer discusses the preconditions and effects of each actions one by one, but rather focuses on the general rules to the domain. These rules apply to not only ‘pick-up’ but also other actions. Therefore, the $\mathbb{DD}$ can be much more concise, requires less algorithmic awareness, and more realistic.
In total, we construct 100 problems varying in complexity for all domains. For each of the two Templated descriptions, there is 1 $\mathbb{DD}$ paired with each of 100 $\mathbb{PD}$s. For the Natural description, there are 100 different pairs of $\mathbb{DD}$s and $\mathbb{PD}$s. We refer to these dataset as BlocksWorld-100, MysteryBlocksWorld-100, Logistics-100 and Barman-100. Data examples can be found in Appendix A.
@@ 4.2 Metrics
Following past work Guan et al. (2023); Zhu et al. (2024), the model-predicted plan is validated using VAL Howey et al. (2004) against the ground-truth $\mathbb{DF}$ and $\mathbb{PF}$ provided above, instead of being compared against “ground-truth” plans like some work Lyu et al. (2023); Liu et al. (2023b); Pan et al. (2023) since there could be multiple correct plans. For the LLM-as-formalizer approach, the predicted $\mathbb{DF}$ and $\mathbb{PF}$ are similarly not compared against the ground-truth, as only the eventual plan is validated because there might be more than one way to formalize the planning domain and problem in PDDL.
We evaluate the predicted plans following Zuo et al. (2024): solvability and correctness. Solvability only applies to LLM-as-formalizer and indicates the percentage solvable predicted $\mathbb{DF}$ and $\mathbb{PF}$, regardless of whether the resulting plan can be executed bsaed on the the gold $\mathbb{DF}$ and $\mathbb{PF}$. Correctness indicates the percentage of actually correct plans. Solvability was determined using the planner dual-bfws-ffparser implemented by Muise (2016) and Correctness was evaluated using VAL333nms.kcl.ac.uk/planning/software/val.html.
Figure 4: Performance of LLM-as-planner (as-P) and LLM-as-formalzier (as-F) across different naturalness level of description on BlocksWorld-100 and Logistics-100. Detailed results are shown in Appendix C.

### v3

@@ Abstract
Large Language Models have been found to create plans that are neither executable nor verifiable in grounded environments. An emerging line of work demonstrates success in using the LLM as a formalizer to generate a formal representation of the planning domain in some language, such as Planning Domain Definition Language (PDDL). This formal representation can be deterministically solved to find a plan. We systematically evaluate this methodology while bridging some major gaps. While previous work only generates a partial PDDL representation, given templated, and therefore unrealistic environment descriptions, we generate the complete representation given descriptions of various naturalness levels. Among an array of observations critical to improve LLMs’ formal planning abilities, we note that most large enough models can effectively formalize descriptions as PDDL, outperforming those directly generating plans, while being robust to lexical perturbation. As the descriptions become more natural-sounding, we observe a decrease in performance and provide detailed error analysis. [^1] Our code and data can be found at https://github.com/CassieHuang22/llm-as-pddl-formalizer.
@@ 1 Introduction
Large Language Models (LLMs) can make informal plans, such as suggesting ideas for parties or giving general advice on immigration. However, most users, let alone automated agents like robots, would not be able to actually execute those plans step-by-step to fruition – either to organize parties or acquire visas – without significant prior knowledge or external help. This inability to make executable plans lies in LLMs’ inability of grounding and formal reasoning Liu et al. (2023b); Pan et al. (2023); Zhang et al. (2023). Cutting-edge research in the community has evaluated LLMs’ ability to make formal plans in grounded environments, such as textual simulations, where all objects and actions represent actualities in the real world. Therefore, any resulting plan that formally involves those objects and actions would be executable and verifiable by nature. Although formal planning has been desirable in the history of AI Weld (1999), recent work found that even state-of-the-art LLMs are unable to generate formal plans Silver et al. (2024); Valmeekam et al. (2024a); Stechly et al. (2024).
@@
Instead of using the LLM as a planner to generate the plan directly, an alternative line of work uses the LLM as a formalizer. Here, the LLM generates a formal representation of a planning domain, for example in PDDL, based on some natural language descriptions of the environment. This representation can then be fed into a solver to find the plan deterministically (see Figure 1). Previous work achieved great success by showing that LLM-as-formalizer greatly outperforms LLM-as-planner in various domains Lyu et al. (2023); Xie et al. (2023); Liu et al. (2023a); Zhang et al. (2024a); Zuo et al. (2024); Zhang et al. (2024c); Zhu et al. (2024), as LLMs are more capable of information extraction than formal reasoning Zhang et al. (2024b). However, the above work has two major shortcomings. First, to simplify the task and evaluation, most have only attempted to generate a partial PDDL representation while assuming the rest is provided, which is often unrealistic in real life. Second, the language used to describe the environments is often artificially templated and structured, leading to potential overestimation of models’ ability.
This paper explores the limit of LLM-as-formalizer devoid of the above two simplifications. We use LLMs to generate the entirety of a PDDL representation, including the domain file and the problem file, given a natural-sounding description of the environment and the task (see Figure 1). On four widely used planning simulations from the International Planning Competition, we benchmark a suite of LLMs on generating PDDL that is both solvable and correct. As the descriptions in these datasets are templated, we also construct model-generated, human-validated descriptions that are natural-sounding to different levels.
Our work is the first to systematically analyze state-of-the-art LLMs’ ability to use the trending methodology of LLM-as-formalizer on the highly challenging task of formal planning. We put forward an array of observations that will benefit future efforts. Discussed in detail in Section 5, our key findings are as follows.
  • Strongest models like gpt-4o, o3-mini, or DeepSeek-R1 can not only plan but also formalize entire PDDL, while other models like Gemma-2 or Llama-3.1 (up to 405B) cannot do either.
  • LLM-as-formalizer often outperforms LLM-as-planner for many model-data combinations, though not always.
  • A more human-like, natural description of the planning environment creates more challenges for models.
  • LLM-as-formalizer is much more robust to long-tail lexical distribution than LLM-as-planner.
  • Weaker models commonly succumb to syntax errors, while all models face semantic errors.
@@ 2 Task: Formal Planning with PDDL
Formal planning (or classical planning) with PDDL involves a domain file ($\mathbb{DF}$) and problem file ($\mathbb{PF}$) (Figure 1). The $\mathbb{DF}$ describes general properties in a planning domain that hold true across problems, while the $\mathbb{PF}$ describes specific configurations of each problem instance. Concretely, the $\mathbb{DF}$ defines all available actions for a state-based environment as well as predicates that represent the properties of object types. Each action definition contains the name of the action, parameters and semantics. The semantics of an action includes the preconditions which describe the necessary world states where the action is valid to execute, and effects which describe how the states change after the action is executed. The $\mathbb{PF}$ defines the involved objects, initial states, and goal states. These two files are then given to a deterministic planner which will algorithmically search for a plan. Such a plan is a series of executable, instantiated actions that sequentially transition the world states from initial to goal. In the AI community, classical planning has been deemed an effective approach to solve real-world users’ problems, as the process is precise, explainable, verifiable, and deterministic.
However, formal planning demands a well-crafted $\mathbb{DF}$ and $\mathbb{PF}$ pair. In a real-world planning scenario, an average user would not describe their environment and problem with PDDL, but more likely with a textual domain description ($\mathbb{DD}$) and problem description ($\mathbb{PD}$). While the $\mathbb{DD}$ and $\mathbb{PD}$ technically describe all the necessary information from the $\mathbb{DF}$ and $\mathbb{PF}$ in natural language, information that can be easily inferred by humans is omitted as the language in the descriptions gets more natural. For example, when describing the action “picking up a block”, a very natural $\mathbb{DD}$ would omit the effect that our hand is no longer empty. These descriptions are then translated into PDDL, which can then be used to find a plan when given to a solver. Hence, we focus on the textual flavor of formal planning: given a $\mathbb{DD}$ and $\mathbb{PD}$, the model outputs a successful plan with regard to the $\mathbb{DF}$ and $\mathbb{PF}$ that are withheld from the model.
@@ 3 Methodology: LLM-as-Formalizer
To tackle the task above, recent work involving LLMs diverged into two methodologies. The first, LLM-as-planner, directly uses LLMs to generate a plan based on the $\mathbb{DD}$ and $\mathbb{PD}$. The second, LLM-as-formalizer, uses LLMs to recover the $\mathbb{DF}$ and $\mathbb{PF}$, before using a deterministic planner to arrive at the plan (Figure 2). Our work will focus on the second while using the first as a baseline. LLM-as-formalizer is in essence neuro-symbolic, where LLMs help define the structured representation that is otherwise prohibitively costly to annotate and brittle to generalize. Existing works in this line demonstrated success but only generate a partial PDDL representation, while assuming the rest. For example, some works generate only the $\mathbb{PF}$ goals while the $\mathbb{DF}$ and remaining $\mathbb{PF}$ are given Lyu et al. (2023); Xie et al. (2023). Others generate the entire $\mathbb{PF}$ while the $\mathbb{DF}$ is given Liu et al. (2023a); Zhang et al. (2024a); Zuo et al. (2024). There have also been works that generate the action semantics in the $\mathbb{DF}$ while the remaining parts of the $\mathbb{DF}$ and entire $\mathbb{PF}$ are provided Zhang et al. (2024c); Zhu et al. (2024), and work that generates the $\mathbb{DF}$ while the $\mathbb{PF}$ is given Wong et al. (2023); Guan et al. (2023). While these assumptions of given PDDL components simplify the task and evaluation, they are often unrealistic. Therefore, we bridge this gap by asking the LLM to predict the entire PDDL – both the $\mathbb{DF}$ and $\mathbb{PF}$ [^2] It is however minimally necessary to provide the action space, the identifiers and parameters of the actions in $\mathbb{DF}$, so the agent knows what actions are possible.
@@ 4 Evaluation: Datasets, Metrics, Models
To evaluate both approaches above, we work with fully-observed textual environments. Here, the provided $\mathbb{DD}$ and $\mathbb{PD}$ contain all necessary information for the model to make a complete plan.
@@ 4.1 Datasets
We consider four simulated planning domains, BlocksWorld, Logistics, Barman from the International Planning Competition IPC (1998), and Mystery BlocksWorld Valmeekam et al. (2024a).
BlocksWorld, also used in Liu et al. (2023a), is a domain to rearrange stacks of blocks on a table using a robotic arm. The domain includes 1 type of entity, 5 predicates, and 4 actions.
Mystery BlocksWorld obfuscates the original BlocksWorld domain by replacing all the names of the types, predicates, actions, and objects with nonsensical words, akin to a wug test Berko (1958). This dataset as a control group can effectively test whether models create plans via lexical pattern-matching and memorization.
@@
Barman, also used in Zhu et al. (2024), is a domain to create cocktails from ingredients using different containers and two robotic arms. In this domain, there are 7 types of entities, 13 predicates, and 12 actions.
Each dataset comes with ground-truth PDDL files describing the domain ($\mathbb{DF}$) and problem ($\mathbb{PF}$). The input to the model is a natural language description of the domain ($\mathbb{DD}$) and problem ($\mathbb{PD}$). The output of the model is a plan, namely a sequence of instantiated actions defined in the $\mathbb{DF}$. For each of these datasets, the natural language description ($\mathbb{DD}$s and $\mathbb{PD}$s) were created in 3 different levels of naturalness.
Heavily Templated. For BlocksWorld, Logistics and Barman, the Heavily Templated $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as Mystery BlocksWorld Valmeekam et al. (2024a). This description is almost a word-by-word translation of PDDL. For example, for the ‘pick-up’ action in BlocksWorld, the ground-truth PDDL $\mathbb{DF}$ would be the following:
@@
From an application point of view, spelling out all preconditions and effects in terms of the predicates is paradoxical, as it assumes the user already has the algorithmic awareness of PDDL.
Moderately Templated. The $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as the original BlocksWorld dataset, following Valmeekam et al. (2024a). For example, the Moderately Templated description of the ‘pick-up’ action is:
@@
While more natural-sounding than the Heavily Templated version, the description still explicitly discusses the preconditions and effects as well as predicates like ‘clear’.
@@
To generate $\mathbb{DD}$s, we ask gpt-4o with high temperature to generate and paraphrase a seed annotated $\mathbb{DD}$, and then manually verify the correctness by making sure it lists the correct predicates, preconditions and effects, which are not unique. We next verify the naturalness of the generated text by making sure there were variations in language throughout all descriptions generated, but unnecessary information was not given out.
To generate $\mathbb{PD}$s, we provide the model with a symbolic configuration containing the initial and goal states of the problem. The model then ‘humanizes’ the problem by making it sound natural, given a couple of seed exemplars. We manually verify the correctness of the datasets of the non-templated problems by hand by comparing them against the problem configurations. We then verify the naturalness of the $\mathbb{PD}$ by making sure there is variation but no ambiguity in its language.
Figure 3: Performance of both usages of LLMs on Heavily Templated BlocksWorld-100, Logistics-100, and MysteryBlocksWorld-100. Detailed results are shown in Appendix D. Due to the cost and zero solvability on our easiest dataset, results of Llama-405B and DeepSeek-8B|70B are omitted.
In the following example of a Natural description of BlocksWorld:
@@
The example no longer discusses the preconditions and effects of each action one by one, but rather focuses on the general rules to the domain. These rules apply to not only ‘pick-up’ but also other actions. Therefore, the $\mathbb{DD}$ can be much more concise, requires less algorithmic awareness, and is more realistic.
In total, we construct 100 problems varying in complexity for all domains. For each of the two Templated descriptions, there is one $\mathbb{DD}$ paired with each of 100 $\mathbb{PD}$s. For the Natural descriptions, there are 100 different pairs of $\mathbb{DD}$s and $\mathbb{PD}$s. We refer to these datasets as BlocksWorld-100, MysteryBlocksWorld-100, Logistics-100 and Barman-100. Moderately Templated and Natural datasets are available only for BlocksWorld-100 and Logistics-100 due to the size of Barman-100. Data examples can be found in Appendix A.
@@ 4.2 Metrics
Following past work Guan et al. (2023); Zhu et al. (2024), the plan produced from LLM-as-Planner is validated using VAL Howey et al. (2004) against the ground-truth $\mathbb{DF}$s and $\mathbb{PF}$s provided above. This is done instead of being compared against “ground-truth” plans like some work Lyu et al. (2023); Liu et al. (2023b); Pan et al. (2023) since there could be multiple correct plans. Similarly, the predicted $\mathbb{DF}$ and $\mathbb{PF}$ for the LLM-as-formalizer approach are not compared against the ground-truth, as only the eventual plan is validated because there might be more than one way to formalize the planning domain and problem in PDDL.
We evaluate the predicted plans following Zuo et al. (2024): using solvability and correctness. Solvability only applies to LLM-as-formalizer and indicates the percentage of solvable predicted $\mathbb{DF}$s and $\mathbb{PF}$s, regardless of whether the resulting plan can be executed based on the ground-truth $\mathbb{DF}$ and $\mathbb{PF}$. Correctness indicates the percentage of actually correct plans. Solvability was determined using the planner dual-bfws-ffparser implemented by Muise (2016) and Correctness was evaluated using VAL333nms.kcl.ac.uk/planning/software/val.html. More information about the solver and validator can be found in the Appendix, Section C.